OpenStudy (anonymous):

Which expression is equal to (F+g)(x) f(x)=x^2+10; g(x)=x-3 a: x^2+x+13 b: x^2+x+7 c: 7x^2+x-3 d: 2x^2+7

jimthompson5910 (jim_thompson5910):

`(f+g)(x)` is the same as `f(x) + g(x)`

jimthompson5910 (jim_thompson5910):

so all you have to do is add the two functions and combine like terms where necessary

OpenStudy (anonymous):

Huh??

jimthompson5910 (jim_thompson5910):

they want you to add `x^2+10` and `x-3`

OpenStudy (anonymous):

Yeah.. And how do i do that?

jimthompson5910 (jim_thompson5910):

do you see any like terms?

OpenStudy (anonymous):

10 and -3

jimthompson5910 (jim_thompson5910):

they add to what?

OpenStudy (anonymous):

7?

jimthompson5910 (jim_thompson5910):

are there any other like terms

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

so we simply say `x^2 + x + 7`

OpenStudy (anonymous):

OHHHHHHH!!!! I get it now!!

OpenStudy (anonymous):

Can you help me with another problem?

jimthompson5910 (jim_thompson5910):

sure, one more

OpenStudy (anonymous):

What is the domain of (f/g)(x) F(X)=(square toot of) 5-x G(X)= (square root of) x+5 a:(xlx(lessthanorequalto)5 b:(xlx(lessthan)5 c:(xl-5<x,(orequalto)5) d:(xlx>(orequaltoo)-5

jimthompson5910 (jim_thompson5910):

solve `x + 5 > 0` to get what?

OpenStudy (anonymous):

IDK

jimthompson5910 (jim_thompson5910):

hint: think of it as `x+5 = 0`

OpenStudy (anonymous):

oh +5 to each side!

jimthompson5910 (jim_thompson5910):

subtract 5 from each side

OpenStudy (anonymous):

oh.. thats what i meant

jimthompson5910 (jim_thompson5910):

so we get x > -5

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

now we solve `5-x >= 0` to get 5-x >= 0 5-x+x >= 0+x 5 >= x x <= 5 so we know that `x > -5` and `x <= 5` put together, we get `-5 < x <= 5`

OpenStudy (anonymous):

Is that the answer?

jimthompson5910 (jim_thompson5910):

yes, it is \(\Large -5 < x \le 5 \)

OpenStudy (anonymous):

Thanks so much!

jimthompson5910 (jim_thompson5910):

np